solve this system by substitution
8x-4y=16 y=2x-4

Answer by benni1013(206)

8x-4y=16
y=2x-4
There is a simpler form of solving this. Do you remeber the following from pre-algebra:
Solve: a-bc when a=2 b=3 and c=1 where you substitute each letter in this case you would get the answer -1. Well instead just numbers we are doing whole equations. In this case we know that y=2x-4. So substitute y=2x-4 into the first equation.
8x-4(2x-4)=16==>8x-8x+16=16==>0x+16=16==>0x=0==>0
Use the distributive property to get to the second step. 8x-8x. Divide by zero to get the answer of 0 for x.
y=2x-4==>2(0)-4==>0-4==>-4
Plug in 0. 0-4. -4.
The answer in ordered pair form is (0,-4)